# Evaluation, Analysis and Mitigation in Multimodal Large Language Models

### v1

@@ Abstract
Hallucination issues persistently plagued current multimodal large language models (MLLMs). While existing research primarily focuses on object-level or attribute-level hallucinations, sidelining the more sophisticated relation hallucinations that necessitate advanced reasoning abilities from MLLMs. Besides, recent benchmarks regarding relation hallucinations lack in-depth evaluation and effective mitigation. Moreover, their datasets are typically derived from a systematic annotation process, which could introduce inherent biases due to the predefined process. To handle the aforementioned challenges, we introduce Reefknot, a comprehensive benchmark specifically targeting relation hallucinations, consisting of over 20,000 samples derived from real-world scenarios. Specifically, we first provide a systematic definition of relation hallucinations, integrating perspectives from perceptive and cognitive domains. Furthermore, we construct the relation-based corpus utilizing the representative scene graph dataset Visual Genome (VG), from which semantic triplets follow real-world distributions. Our comparative evaluation across three distinct tasks revealed a substantial shortcoming in the capabilities of current MLLMs to mitigate relation hallucinations. Finally, we advance a novel confidence-based mitigation strategy tailored to tackle the relation hallucinations problem. Across three datasets, including Reefknot, we observed an average reduction of 9.75% in the hallucination rate. We believe our paper sheds valuable insights into achieving trustworthy multimodal intelligence. Our dataset and code will be released upon paper acceptance.
@@ 1 Introduction
POPE (Li et al. 2023c) COCO Post-processed Acc. Co-occur.
HaELM (Wang et al. 2023b) MS-COCO Post-processed Acc. Attention
MME (Fu et al. 2024) Self-Sourced Manual Acc. -
AMBER (Wang et al. 2024a) MS-COCO Post-processed Acc. -
MHaluBench (Chen et al. 2024b) Self-Sourced Post-processed Prec. -
R-Bench (Wu et al. 2024) COCO Post-processed Acc. Co-occur.
MMRel (Nie et al. 2024) Visual-Genome Synthetic Acc. -
VALOR-EVAL (Qiu et al. 2024) GQA Post-processed LLM-based Co-occur.
@@
In recent years, large language models (LLMs) have revolutionized the AI field by expanding their training data to trillions of tokens and increasing their parameter counts to hundreds of billions (Brown et al. 2020; Achiam et al. 2023; Touvron et al. 2023). This has unlocked powerful emergent abilities, and seen widespread applications in diverse domains (Achiam et al. 2023; Yan et al. 2024; Wang et al. 2023a). Recently, the community managed to combine visual backbones with powerful LLMs, resulting in multimodal large language models (MLLMs) (Liu et al. 2023b). While this has led to advancements in multimodal scenarios, it also presents challenges for MLLMs, notably their tendency to generate hallucinations. In LLMs, hallucinations occur when the model generates inaccurate or misleading factual information that cannot be supported by existing knowledge (Zhang et al. 2023). However, such issues become more complex in MLLMs, as hallucinations can manifest as responses containing references or descriptions of the input image that are incorrect (Bai et al. 2024; Huo et al. 2024a). Therefore, it is crucial to evaluate and mitigate these hallucinations to improve the trustworthy of MLLMs in real-world scenarios.
@@
Figure 1: Comparison between the focus of Reefknot benchmark — relation hallucination with two categories (i.e., perceptive and cognitive) vs. object and attribute hallucinations.
@@
Hallucination in MLLMs likely originates from knowledge biases between pre-training and fine-tuning, including statistical biases in the training data, over-reliance on parametric knowledge, and skewed representation learning, as suggested by previous research (Bai et al. 2024; Zhu et al. 2024). Specifically, the hallucination in MLLMs can be divided into three categories: object, attribute and relation hallucinations (Bai et al. 2024; Liu et al. 2024).
@@
As depicted in Figure 1 (a), object-level hallucination focuses on the model’s discrimination of the existence of basic objects (Li et al. 2023c); while as shown in Figure 1 (b), attribute-level hallucination often focuses on whether the model can distinguish some properties of the object itself like color, number, shape and so on (Fu et al. 2024). Their commonality is that they only focus on the single object present in the image. There has been much work exploring the alleviation of these two types of hallucinations. For example, Woodpecker (Yin et al. 2023) used a post-processing method to correct hallucination after the generation process; VCD (Leng et al. 2023) proposed a contrast decoding strategy to mitigate hallucination by adding noise to images.
@@
Figure 2: The hallucination rates on POPE (Li et al. 2023c), an object hallucination benchmark, and our Reefknot with a focus on relation hallucination (w/ same configuration).
@@
Despite these efforts, the community barely considers relation hallucination, which demands more complex reasoning capabilities from MLLMs. Figure 1 (c) reveals that relation hallucination is related to at least two objects simultaneously shown in the image, through either perceptive or cognitive perspective. Furthermore, as outlined in Table 1, recent highly relevant benchmarks on hallucinations lack thorough evaluation and effective mitigation strategies.
@@
To handle the aforementioned research gaps, we propose the first comprehensive benchmark Reefknot to evaluate the performance on relation hallucination. Unlike many benchmarks of MLLM that were constructed by automatic labeling technique, we construct the relation-based dataset based on semantic triplets retrieved from the scene graph dataset. Table 1 also demonstrates that our triplets are from real-life scenarios, without any post-processing (e.g., segmentation and bounding box techniques), manual annotation, and synthetic method (e.g., diffusion-based generation). We categorize relation hallucinations into two types: perceptive, involving concrete relational terms like ”on”, ”in”, ”behind” and cognitive, which encompasses more abstract terms such as ”blowing” and ”watching”. Second, we evaluate the mainstream MLLMs on Reefknot via three diverse tasks across two types of relation hallucinations. Figure 2 also illustrates that relation hallucination can be more severe than object hallucination in current MLLMs, highlighting the importance of our evaluation. Furthermore, we propose a simple relation hallucination mitigation method named Detect-then-Calibrate. This originates from the experimental observation that when relation hallucinations occur, the response probability drops significantly, hovering just above 50% in extreme cases compared to the usual nearly 90%. Our method achieves an average improvement of 9.75% across three relation hallucination benchmarks.
In summary, Our main contributions are as follows:
@@ 2 Relation Hallucination Benchmark
In this section, we describe the dataset construction pipeline of Reefknot benchmark as shown in Figure 3. Unlike object and attribute hallucinations that only involve one entity, relations involve three entities making it more difficult to handle. We first identify relation triplets from Visual Genome (VG) dataset (Krishna et al. 2016) (Phase a), and conduct triplet filtering (Phase b). Subsequently, we extract the semantic triplets (Phase c) and categorize their relations (Phase d). Then, a relation-based question set can be constructed into three types (Phase e). Finally, the quality of dataset is ensured by three rounds of expert-based validation (Phase f).
@@
The dataset comprises of 11,084 images taken from VG dataset, a finely annotated scene graph dataset utilized by the research community (Tang et al. 2020; Liang et al. 2019). As indicated in Figure 3 (a), visual objects and their relations from VG dataset can be easily identified. Besides, we filter the triplets with redundant information, incorrect relationships, or noisy descriptions, as depicted in Figure 3 (b). Subsequently, we can extract semantic triplets by identifying subject-object pairs and the relationships between them, forming (subject, relation, object) triplets in Figure 3 (c).
@@
As depicted in Figure 3 (d), we categorize relationships into two categories based on deep semantic meanings: perceptive and cognitive. Perceptive relationships involve locational prepositions, such as <boy, behind, sofa>and <cup, on, table>; whereas cognitive relationships are expressed through action phrases indicating states, such as <boy, eating, food>and <girl, sleeping in, bed>. ChatGPT is employed to assist in this classification. Table 2 also indicates the sample numbers in different tasks and hallucination categories. The prompt we used for relation categorization is decipted in Figure 4.
Prompt: ————————————————– You are a relation term classification assistant. Please help me determine whether the following relational terms/phrases belong to perceptive relationships or cognitive relationships. Perceptive relationships are defined as those involving locational or state-based prepositions, such as “on” and “behind”. Cognitive relationships are defined as some words that indicate an action or behavior, such as holding, watching, etc. Here I will give you some demonstrations for reference: {In-context examples} Please answer with a single word in <Perception, Cognition >. Input: {Input}
Figure 4: Prompt template for relation categorization.
@@
Table 2: Statistical overview of Reefknot benchmark dataset.
@@
Phi-3-vision-128k-instruct (Abdin et al. 2024) 4.2B 39.88 57.07 50.98 33.97 21.35 49.45 60.30
Yi-VL (AI et al. 2024) 6B 33.56 47.53 71.02 33.16 16.33 74.96 55.81
LLaVA (Liu et al. 2023a) 7B 37.67 68.05 52.93 33.99 51.04 54.56 51.41
MiniGPT4-v2 (Chen et al. 2023) 7B 46.7 78.00 61.30 43.73 68.50 65.88 39.88
MiniCPM (Hu et al. 2024) 7B 31.93 48.65 47.63 27.65 16.71 45.96 65.73
Qwen-VL (Bai et al. 2023) 7B 42.21 56.7 72.47 33.53 21.88 73.01 52.55
Deepseek-VL (DeepSeek-AI et al. 2024) 7B 37.58 56.33 67.07 32.22 23.60 59.34 56.39
GLM4V (GLM et al. 2024) 9B 34.09 50.47 58.09 27.08 16.87 56.47 62.03
LLaVA (Liu et al. 2023a) 13B 40.7 59.35 48.93 34.19 29.19 54.45 57.47
CogVLM (Wang et al. 2024b) 19B 37.23 47.95 70.14 29.89 18.54 66.18 57.1
Yi-VL (AI et al. 2024) 34B 32.79 44.19 57.67 33.75 14.85 52.72 62.61
GPT-4o (OpenAI et al. 2024) - 32.56 40.93 42.70 26.27 11.53 48.78 68.32
@@
  • For Yes/No (Y/N) questions, we employ an adversarial approach by introducing a negative sample within the same triplet, alongside the positive sample, to test the model’s ability to correctly answer “No”.
  • Multiple Choice Questions (MCQ) are designed with one correct answer and three random options to evaluate the model’s resistance to relation hallucinations within a controlled vocabulary.
@@
Finally, we perform multi-turn manual verification to ensure the quality of the question sets (see Figure 3 (f)). Each question undergoes at least three rounds of review by four domain experts. We revise any inappropriate expressions and exclude meaningless questions, such as “Is the window on the wall?”, which lack informative value and can be answered without visual input. After rigorous screening process, our dataset comprises 21,880 questions across 11,084 images as shwon in Table 2.
@@ 3.1 Task and Evaluation Metrics
As illustrated in Table 3, we conducted evaluations on mainstream MLLMs to evaluate relation-level hallucinations. All experiments of open-sourced MLLMs were conducted using 8 NVIDIA A100 GPUs. To reduce variability, each experiment was run three times, and we reported the average of these results. In our evaluation, we reported our results from two distinct categories: perception and cognition.
For discriminative Y/N and MCQ tasks, we reported the hallucination rate $Halr$ as a metric. Generative questions have always posed a challenge in evaluation of MLLMs. In our assessment of the generative VQA task, we employed the DeBERTa model using a bidirectional entailment approach for label match (Kuhn, Gal, and Farquhar 2023), and we denoted $Halr$ as metric for simplicity as well. For more details, we have listed in Appendix. In general, we use the following metric $R_{\text{score}}$ to comprehensively evaluate the overall performance across the three tasks:
@@ 3.2 Main results
Significant performance differences among the various models are evident. Table 3 shows significant differences in the performance of various models across different question types. For instance, Qwen-vl-chat excels in Y/N and MCQ settings but encounters serious hallucination issues in VQA tasks. A detailed review of the model’s responses reveals that while Qwen can follow instructions accurately, many responses contain expressions that are completely unrelated to the labels. We presume that such a phenomenon is owing to over-fitting training during the instruction-tuning process. Among open-sourced models, MiniCPM (Hu et al. 2024) stands out, likely due to its adoption of the fine-grained alignment technique such as RLHF-V (Yu et al. 2023) to alleviate hallucinations.
Figure 6: Confusion matrixes of MiniCPM-7B on Reefknot benchmark (Left: Y/N setting; Right: MCQ setting).
@@
Figure 5 presents a comparative analysis of the top five best models’ performances on the Reefknot benchmark. The comparison reveals a clear hierarchy in performance, with GPT-4o outperforming the rest across all tasks. MiniCPM, Yi-VL-34B, and GLM4v-9B show a competitive edge, closely following GPT-4o, while Phi-3 lags behind relatively. In terms of task performance, Figure 5 indicates that while some models excel in one task, they may not perform as well in others. For instance, Yi-VL-34B has the second-highest score in MCQ but falls short in VQA. This highlights the importance of task-specific model tuning and the need for a balanced approach to improve overall performance across all relation hallucination tasks.
@@
To analyze the error cases quantitatively, we visualized the results of MiniCPM-7B, the best open-sourced model, across two discriminative settings, as illustrated in Figure 6. For Y/N questions, the model tends to favor positive responses (i.e., Yes). Specifically, among all misclassifications, the instances where a No label is incorrectly classified as Yes are twice as times as instances where a Yes label is incorrectly classified as No. Besides, the model tends to answer D in MCQ settings. We suspect the aforementioned tendency is likely due to the imbalance in the distribution of the training data.

### v2

# Hallucination Evaluation, Analysis and Mitigation in Multimodal Large Language Models

@@ Abstract
Hallucination issues continue to affect multimodal large language models (MLLMs), with existing research mainly addressing object-level or attribute-level hallucinations, neglecting the more complex relation hallucinations that require advanced reasoning. Current benchmarks for relation hallucinations lack detailed evaluation and effective mitigation, and their datasets often suffer from biases due to systematic annotation processes. To address these challenges, we introduce Reefknot, a comprehensive benchmark targeting relation hallucinations, comprising over 20,000 real-world samples. We provide a systematic definition of relation hallucinations, integrating perceptive and cognitive perspectives, and construct a relation-based corpus using the Visual Genome scene graph dataset. Our comparative evaluation reveals significant limitations in current MLLMs’ ability to handle relation hallucinations. Additionally, we propose a novel confidence-based mitigation strategy, which reduces the hallucination rate by an average of 9.75% across three datasets, including Reefknot. Our work offers valuable insights for achieving trustworthy multimodal intelligence. The dataset and code are released at https://github.com/JackChen-seu/Reefknot.
Reefknot: A Comprehensive Benchmark for Relation
Hallucination Evaluation, Analysis and Mitigation in Multimodal Large Language Models
@@ 1 Introduction
POPE Li et al. (2023c) COCO Post-processed Acc. Co-occur.
HaELM Wang et al. (2023b) MS-COCO Post-processed Acc. Attention
MME Fu et al. (2024) Self-Sourced Manual Acc. -
AMBER Wang et al. (2024a) MS-COCO Post-processed Acc. -
MHaluBench Chen et al. (2024b) Self-Sourced Post-processed Prec. -
R-Bench Wu et al. (2024) COCO Post-processed Acc. Co-occur.
MMRel Nie et al. (2024) Visual-Genome Multi-source Acc. -
VALOR-EVAL Qiu et al. (2024) GQA Post-processed LLM-based Co-occur.
@@
In recent years, large language models (LLMs) have revolutionized the AI field by expanding their training data to trillions of tokens and increasing their parameter counts to hundreds of billions Brown et al. (2020); Achiam et al. (2023); Touvron et al. (2023). This has unlocked powerful emergent abilities, and seen widespread applications in diverse domains Achiam et al. (2023); Yan et al. (2024); Wang et al. (2023a). Recently, the community managed to combine visual backbones with powerful LLMs, resulting in multimodal large language models (MLLMs) Liu et al. (2023b). While this has led to advancements in multimodal scenarios, it also presents challenges for MLLMs, notably their tendency to generate hallucinations Zou et al. (2024). In LLMs, hallucinations occur when the model generates inaccurate or misleading factual information that cannot be supported by existing knowledge Zhang et al. (2023). However, such issues become more complex in MLLMs, as hallucinations can manifest as responses containing references or descriptions of the input image that are incorrect Bai et al. (2024); Huo et al. (2024a). Therefore, it is crucial to evaluate and mitigate these hallucinations to improve the trustworthy of MLLMs in real-world scenarios.
@@
Figure 1: Comparison between the focus of Reefknot — relation hallucination with two categories (i.e., perceptive & cognitive) vs. object & attribute hallucinations.
@@
Hallucination in MLLMs likely originates from knowledge biases between pre-training and fine-tuning, including statistical biases in the training data, over-reliance on parametric knowledge, and skewed representation learning, as suggested by previous research Bai et al. (2024); Zhu et al. (2024). Specifically, the hallucination in MLLMs can be divided into three categories: object, attribute and relation hallucinations Bai et al. (2024).
@@
As depicted in Figure 1 (a), object-level hallucination focuses on the model’s discrimination of the existence of basic objects Li et al. (2023c); while as shown in Figure 1 (b), attribute-level hallucination often focuses on whether the model can distinguish some properties of the object itself like color, number, shape and so on Fu et al. (2024). Their commonality is that they only focus on the single object present in the image. There has been much work exploring the alleviation of these two types of hallucinations. For example, Woodpecker Yin et al. (2023) used a post-processing method to correct hallucination after the generation process; VCD Leng et al. (2023) proposed a contrast decoding strategy to mitigate object-level hallucination by adding noise to images.
@@
Figure 2: The hallucination rates on POPE, an object hallucination benchmark, and our Reefknot with a focus on relation hallucination (w/ same configuration).
@@
Despite these efforts, the community barely considers relation hallucination, which demands more complex reasoning capabilities from MLLMs. Figure 1 (c) reveals that relation hallucination is related to at least two objects simultaneously in the image, through either perceptive or cognitive perspective. Furthermore, as outlined in Table 1, recent highly relevant benchmarks on hallucinations lack thorough evaluation and effective mitigation strategies.
@@
To handle the aforementioned research gaps, we propose the first comprehensive benchmark Reefknot to evaluate the performance on relation hallucination. Unlike many benchmarks of MLLM that were constructed by automatic labeling technique, we construct the relation-based dataset based on semantic triplets retrieved from the scene graph dataset. Table 1 demonstrates that our triplets are from real-life scenarios, without any post-processing (e.g., segmentation and bounding box techniques), manual annotation, and synthetic method (e.g., diffusion-based generation). We categorize relation hallucinations into two types: perceptive, involving concrete relational terms like "on", "in", "behind" and cognitive, which encompasses more abstract terms such as "blowing" and "watching". Second, we evaluate the mainstream MLLMs on Reefknot via three diverse tasks across two types of relation hallucinations. Figure 2 also illustrates that relation hallucination can be more severe than object hallucination in current MLLMs, highlighting the importance of our evaluation. Furthermore, we propose a simple relation hallucination mitigation method named Detect-then-Calibrate. This originates from the experimental observation that when relation hallucinations occur, the response probability drops significantly, hovering just above 50% in extreme cases compared to the usual nearly 90%. Our method achieves an average improvement of 9.75% across three relation hallucination benchmarks. In summary, Our main contributions are as follows:
@@
    We investigated the mechanism of relation hallucination generation from the perspective of response confidence and identified a correlation between relation hallucination and high uncertainty from token levels.
  4. 4.
@@ 2 Relation Hallucination Benchmark
In this section, we describe the dataset construction pipeline of Reefknot benchmark in Figure 3. Unlike object and attribute hallucinations that only involve one entity, relations involve two entities making it more difficult to handle. We first identify relation triplets from Visual Genome (VG) dataset Krishna et al. (2016) (Phase a), and conduct triplet filtering (Phase b). Subsequently, we extract the semantic triplets (Phase c) and categorize their relations (Phase d). Then, a relation-based question set can be constructed into three types (Phase e). Finally, the quality of dataset is ensured by three rounds of expert-based validation (Phase f).
@@
The dataset comprises of 11,084 images taken from VG dataset, a finely annotated scene graph dataset utilized by the research community Tang et al. (2020); Liang et al. (2019). As indicated in Figure 3 (a), visual objects and their relations from VG dataset can be easily identified. Besides, we filter the triplets with redundant information, incorrect relationships, or noisy descriptions, as depicted in Figure 3 (b). Subsequently, we can extract semantic triplets by identifying subject-object pairs and the relationships between them, forming (subject, relation, object) triplets in Figure 3 (c).
@@
As depicted in Figure 3 (d), we categorize relationships into two categories based on deep semantic meanings: perceptive and cognitive. Perceptive relationships involve locational prepositions, such as <boy, behind, sofa >and <cup, on, table>; whereas cognitive relationships are expressed through action phrases indicating states, such as <boy, eating, food>and <girl, sleeping in, bed>. ChatGPT is employed to assist in this classification. Table 2 also indicates the sample numbers in different tasks and hallucination categories. The prompt we used for relation categorization is listed in Appendix A.4.
@@
Table 2: Detailed statistical overview of the Reefknot benchmark dataset, including sample distribution across categories and task types.
@@
Phi-3-vision-128k-instruct Abdin et al. (2024) 4.2B 39.88 57.07 50.98 33.97 21.35 49.45 60.30
Yi-VL AI et al. (2024) 6B 33.56 47.53 71.02 33.16 16.33 74.96 55.81
LLaVA Liu et al. (2023a) 7B 37.67 68.05 52.93 33.99 51.04 54.56 51.41
MiniGPT4-v2 Chen et al. (2023) 7B 46.7 78.00 61.30 43.73 68.50 65.88 39.88
MiniCPM Hu et al. (2024) 7B 31.93 48.65 47.63 27.65 16.71 45.96 65.73
Qwen-VL Bai et al. (2023) 7B 42.21 56.7 72.47 33.53 21.88 73.01 52.55
Deepseek-VL DeepSeek-AI et al. (2024) 7B 37.58 56.33 67.07 32.22 23.60 59.34 56.39
GLM4V GLM et al. (2024) 9B 34.09 50.47 58.09 27.08 16.87 56.47 62.03
LLaVA Liu et al. (2023a) 13B 40.7 59.35 48.93 34.19 29.19 54.45 57.47
CogVLM Wang et al. (2024b) 19B 37.23 47.95 70.14 29.89 18.54 66.18 57.1
Yi-VL AI et al. (2024) 34B 32.79 44.19 57.67 33.75 14.85 52.72 62.61
GPT-4o OpenAI et al. (2024) - 32.56 40.93 42.70 26.27 11.53 48.78 68.32
@@
  • For Yes/No (Y/N) questions, we employ an adversarial approach by introducing a negative sample within the same triplet, alongside the positive sample, to test the model’s ability to correctly answer "No".
  • Multiple Choice Questions (MCQ) are designed with one correct answer and three random options to evaluate the model’s resistance to relation hallucinations within a controlled and limited vocabulary.
@@
Finally, we perform multi-turn manual verification to ensure the quality of the question sets (see Figure 3 (f)). Each question undergoes at least three rounds of review by four domain experts. We revise any inappropriate expressions and exclude meaningless questions, such as “Is the window on the wall?", which lack informative value and can be answered without visual input. After rigorous screening process, our dataset comprises 21,880 questions across 11,084 images as shwon in Table 2.
@@ 3.1 Task and Evaluation Metrics
As illustrated in Table 3, we conducted evaluations on mainstream MLLMs to evaluate relation-level hallucinations. All experiments of open-sourced MLLMs were conducted using 8 NVIDIA A100 GPUs. Each experiment ran three times, and we reported the average of these results. In our evaluation, we reported our results from two distinct categories: perception and cognition.
For discriminative Y/N and MCQ tasks, we reported the hallucination rate $Halr$ as a metric. Generative questions have always posed a challenge in evaluation of MLLMs. In our assessment of the generative VQA task, we employed the DeBERTa model using a bidirectional entailment approach for label match Kuhn et al. (2023), and we denoted $Halr$ as metric for simplicity as well. For more details, we have listed in Appendix A.6. In general, we use the following metric $R_{\text{score}}$ to comprehensively evaluate the overall performance across the three tasks, which is expressed in Formula 1.
@@ 3.2 Main results
Significant performance differences among the various models are evident. Table 3 shows significant differences in the performance of various models across different question types. For instance, Qwen-vl-chat excels in Y/N and MCQ settings but encounters serious hallucination issues in VQA tasks. A detailed review of the model’s responses reveals that while Qwen can follow instructions accurately, many responses contain expressions that are completely unrelated to the labels. We presume that such a phenomenon is owing to over-fitting training during the instruction-tuning process. Among open-sourced models, MiniCPM Hu et al. (2024) stands out, likely due to its adoption of the fine-grained alignment technique such as RLHF-V Yu et al. (2023) to alleviate hallucinations.
Figure 5: Confusion matrices of MiniCPM-7B’s performance on Reefknot benchmark (Left: Y/N setting; Right: MCQ setting).
@@
Figure 4 presents a comparative analysis of the top five best models’ performances on the Reefknot benchmark. The comparison reveals a clear hierarchy in performance, with GPT-4o outperforming the rest across all tasks. MiniCPM, Yi-VL-34B, and GLM4v-9B show a competitive edge, closely following GPT-4o, while Phi-3 lags behind relatively. In terms of task performance, Figure 4 indicates that while some models excel in one task, they may not perform as well in others. For example, Yi-VL-34B achieves the second-highest score in MCQ but underperforms in VQA. This underscores the significance of task-specific model tuning and the necessity of a balanced approach to enhance overall performance across all relation hallucination tasks.
@@
To analyze the error cases quantitatively, we visualized the results of MiniCPM-7B, the best open-sourced model, across two discriminative settings, as illustrated in Figure 5. For Y/N questions, the model tends to favor positive responses (i.e., Yes). Specifically, among all misclassifications, the instances where a No label is incorrectly classified as Yes are twice as times as instances where a Yes label is incorrectly classified as No. Besides, the model tends to answer D in MCQ settings. We suspect the tendency is likely due to the imbalance in the distribution of the training data.